a body thrown vertically up with initial velocity 52 m/s from the ground passes twice a point at h height above at an interval of 10s. The height h is...........
(g=10m/s^2)

Grade:

2 Answers

Arun Kumar IIT Delhi
9 years ago
time of flight 2 vv/g=10
vf=50
=>s=(52^2-50^2)/20

10.2m

Khushi singh
6 years ago
As given u=52m/s Aplly s=ut+1/2at2s=52×10-1/2×10×10×10s=520-500s=20As body passes the point twice it is the height if bth the points together so to find the height of one pointh=s/2h= 10m